### v1

@@ 4.2 Conditioning normalizing flows on temperature and pressure
where the importance weights are defined similarly to Eq. (8) and $\Delta f_{AB}$ is the free energy difference between $A$ and $B$. Since $\Delta f_{AB}$ is independent of the trainable weights, the loss function simplifies to
$\mathcal{L}_{AB}(\theta)=-\mathbb{E}_{{\bf x}\sim q_{A}}\bigl{[}\log w({\bf x})\bigr{]}\geq\Delta f_{AB}.$ (13)
For efficient training, the determinant of the Jacobian of the transformation must be easy to calculate (see Eq. (8)). For this purpose, coupling layers [43, 44] were designed which split the input data into two channels and transform one of the channels conditioned on the other. This yields a triangular Jacobian whose determinant can be evaluated analytically. The entire input vector can be transformed by stacking several coupling layers, whereby the channel to be transformed changes.
A conditional normalizing flow $f({\bf x}|{\bf c})$ can be trained to learn a mapping between a prior distribution $q_{A}({\bf x})$ and a conditional target distribution $q_{B}({\bf x}^{\prime}|{\bf c})$ for a vector of conditioning variables ${\bf c}$ [25, 26]. In this case, the change of variable reads
@@
For energy-based training, the trainable parts of the conditional KL divergence are given by [25, 27]
@@
with the conditional importance weights defined as $w({\bf x}|{\bf c})\!=\!q_{B}({\bf x}|{\bf c})/q^{\prime}_{A}({\bf x}|{\bf c})$. In Eq. (15), we used the single-point loss from Eq. (13) and $p_{\bf c}$ defines the distribution of the conditional variable. In practice, conditioning a coupling flow architecture can be easily achieved by concatenating the conditioning variable to the unchanged part before transforming the other part [25].
For the application to free energy sampling within TFEP in the isobaric-isothermal ($NPT$) ensemble, we aim to train a normalizing flow to learn a mapping conditioned on temperature $T$ and pressure $P$ (${\bf c}\!=\!(T,P)$) between equilibrium distributions of the form (see Eq. (3))
$q({\bf x},V|{\bf c})\propto e^{-u({\bf x},V|{\bf c})}.$ (16)
For brevity, we use $q_{\bf c}=q({\bf x},V|{\bf c})$ in the following. In this notation, we want to learn a mapping between a prior distribution $q_{{\bf c}_{0}}=q_{{\bf c}={\bf c}_{0}}$, sampled with MD, and the conditional target distribution $q_{\bf c}$. For training the flow, we minimize the conditional loss function (Eq. (15)) using the importance weights defined in Eq. (10):
$\mathcal{L}(\theta)$ $=-\mathbb{E}_{{\bf c}\sim p_{\bf c}}\mathbb{E}_{({\bf x},V)\sim q_{{\bf c}_{0}}}\bigl{[}\log w({\bf x},V|{\bf c})\bigr{]}$ (17)
@@
which results in the loss function Eq. (1). Once the flow is trained, the free energy difference between the prior distribution and the target distribution is obtained from TFEP (Eq. (9))
@@ 4.3 Flow architecture
Figure 4: Illustration of one $(T,P,V^{\prime})$- conditional coupling layer over atomic coordinates. The sets of fractional coordinates and spline parameters are denoted as $\{s\}=\{s_{1},\dots,s_{N}\}$ and $\{\psi\}=\{\psi_{1},\dots,\psi_{N}\}$, respectively. Dashed lines denote splitting / merging operations.
As sketched in Fig. 1, our flow is based on a shape-conditional architecture [21]. Here, the volume is updated by an affine transformation, while the atomic configurations ${\bf x}\in\mathbb{R}^{3N}$ are first scaled to fractional coordinates ${\bf s}\in[0,1]^{3N}$ and then updated conditioned on the volume and the thermodynamic state.
Our flow transforms a sample of the prior distribution $({\bf x},V)$ to $({\bf x}^{\prime},V^{\prime})$, where the volume is connected to the box dimensions ${\bf L}$ by $V\!=\!\Pi_{i=1}^{3}L_{i}$. In this work, we focus on the simulation of solids and liquids and therefore apply certain restrictions regarding the shape of the simulation box. We use orthorhombic simulation boxes with ${\bf L}=[L_{1},L_{2},L_{3}]$ and further only allow for an isotropic scaling of the box ($L_{i}/L_{j}\!=\!{\rm const.}$), such that the box is fully characterized by its volume $V\!=\!\Pi_{i=1}^{3}L_{i}$.
The volume transformation and the transformation of the position of particle $i$ can be written as
@@ 4.4 Conditional coupling layer
To be applicable to condensed phase system, $f_{\bf s}$ must respect periodic boundary conditions and invariance under particle permutation. In addition, the center of mass needs to be fixed to prevent uncontrolled errors in the free energy prediction due to rigid translations [20].
To achieve these requirements, we extend the coupling layer originally proposed in [18] and refined in [19] to the case of conditional sampling. Figure 4 shows our implementation of a $(T,P,V)$ conditional coupling layer. The $N$-particle configuration ${\bf s}$ is split over cartesian coordinates into two subsets $\{s^{\nu}\}\!=\!\{s^{\nu}_{1},...,s^{\nu}_{N}\}$ and $\{s^{\mu}\}\!=\!\{s^{\mu}_{1},...,s^{\mu}_{N}\}$, e.g. $\nu=x,\mu=\{y,z\}$. First, all coordinates indexed by $\nu$ remain unchanged while the coordinates of atom $i$ carrying index $\mu$ are transformed according to
@@
Here, $G$ is a circular spline [30], which naturally accounts for the periodicity of the problem. The parameters of the spline $\psi_{i}^{\mu}$ are obtained as the output of a transformer [31] $C$ for the coordinates indexed by $\mu$ of particle $i$, which takes as input circular encodings of the coordinates $\nu$ and is conditioned on $T$,$P$, and $V^{\prime}$. The partitioning of $\nu$ and $\mu$ is changed between layers such that each coordinate gets transformed (see [18, 19] for details).
@@
4.5 Computing coexistence lines
The coexistence line in the $(T,P)$ diagram of two phases of a system is defined by equal total Gibbs free energies $\tilde{G}(T,P)$ including the kinetic contribution. For a given state ($T_{i},P_{i}$), $\tilde{G}_{i}$ is the sum of the configurational contribution (Eq. (5)) and the kinetic contribution:
$\tilde{G}_{i}=G_{i}+\frac{3N}{\beta_{\rm i}}\log\Lambda(T_{\rm i}).$ (25)
Here, $\Lambda(T_{i})$ is the thermal De-Broglie wavelength $\Lambda(T_{i})=(\frac{h^{2}}{2\pi mk_{B}T_{i}})^{1/2}$.
For each phase, the configurational contribution $G$ can be obtained by combining relative differences within phases with the absolute free energy $G_{\rm ref}$ at a reference state $(T_{\rm ref},P_{\rm ref}$.) Using the difference in the dimensionless free energy between the reference point and a state $i$, $\Delta f_{{\rm ref},i}$, and writing it explicitly as
$\Delta f_{{\rm ref},i}=\beta_{i}G_{i}-\beta_{\rm ref}G_{\rm ref},$ (26)
the absolute Gibbs free energy can be obtained as
The differences $\Delta f_{{\rm ref},i}$ can be obtained from free energy estimators such as MBAR and TFEP (Eq. (9)) after performing simulations of the physical system.
$G_{\rm ref}=F^{0}_{\rm ref}+\Delta F^{0}_{\rm ref}+P_{\rm ref}V_{\rm ref}$ (28)
In principle, a flow can be used to determine $\Delta f_{{\rm ref},i}$ as well as $\Delta F^{0}_{\rm ref}$. However, it has been shown that obtaining $\Delta F^{0}_{\rm ref}$ is computationally much more expensive than an MD+MBAR based approach [19]. For a flow-based calculation of the coexistence line, we therefore decide to determine $\Delta F^{0}_{\rm ref}$ using MD+MBAR and to use our conditional flow approach to obtain $\Delta f_{{\rm ref},i}$.

### v2

@@ 2.2 Normalizing flows
where the importance weights are defined similarly to Eq. (6) as $w({\bf x})\!=\!q_{B}(f({\bf x}))/q^{\prime}_{A}(f({\bf x}))$ and $\Delta f_{AB}$ is the free energy difference between $A$ and $B$. The loss function simplifies to
$\mathcal{L}_{AB}(\theta)=-\mathbb{E}_{{\bf x}\sim q_{A}}\bigl{[}\log w({\bf x})\bigr{]}\geq\Delta f_{AB},$ (11)
since $\Delta f_{AB}$ is independent of the trainable weights.
For efficient training of normalizing flows, the determinant of the Jacobian of the transformation must be easy to calculate (see Eq. (6)). For this purpose, coupling layers [29, 30] were designed which split the input data into two channels and transform one of the channels conditioned on the other. This yields a triangular Jacobian whose determinant can be evaluated analytically. The entire input vector can be transformed by stacking several coupling layers, whereby the channel to be transformed changes.
Flows can further be trained in a conditional setting which allows the modeling of conditional target distributions [31, 32]. A conditional normalizing flow $f({\bf x}|{\bf c})$ is trained to learn a mapping between a prior distribution $q_{A}({\bf x})$ and a conditional target distribution $q_{B}({\bf x}^{\prime}|{\bf c})$ for a vector of conditioning variables ${\bf c}$ [31, 32]. In this case, the change of variable reads
@@
For energy-based training, the trainable parts of the conditional KL divergence are given by the average of the KL divergences between the generated and the target distributions conditioned on each target state over the entire range of ${\bf c}$, yielding the following loss function:
@@
with the conditional importance weights defined as $w({\bf x}|{\bf c})\!=\!q_{B}({\bf x}|{\bf c})/q^{\prime}_{A}({\bf x}|{\bf c})$. In practice, conditioning a coupling flow architecture can be easily achieved by concatenating the conditioning variable to the unchanged part before transforming the other part [31]. While originally developed in the context of image generation tasks [31, 32], the idea of conditioning flows has recently been applied to the sampling of rare events and to lattice field theories [33, 34].
2.3 Computing coexistence lines
The coexistence line in the $(T,P)$ diagram of two phases of a system is defined by equal total Gibbs free energies $\tilde{G}(T,P)$. For a given state ($T_{i},P_{i}$), $\tilde{G}_{i}$ is the sum of the configurational contribution (Eq. (3)) and the kinetic contribution:
$\tilde{G}_{i}=G_{i}+\frac{3N}{\beta_{\rm i}}\log\Lambda(T_{\rm i}),$ (14)
where $\Lambda(T_{i})$ is the thermal De-Broglie wavelength $\Lambda(T_{i})=(\frac{h^{2}}{2\pi mk_{B}T_{i}})^{1/2}$.
Using free energy estimators such as MBAR or TFEP, the configurational contribution $G$ for each phase can be obtained by combining relative differences within phases with the absolute free energy $G_{\rm ref}$ at a reference state $(T_{\rm ref},P_{\rm ref}$.) Writing the difference in the dimensionless free energy between the reference point and a state $i$, $\Delta f_{{\rm ref},i}$ explicitly as
$\Delta f_{{\rm ref},i}=\beta_{i}G_{i}-\beta_{\rm ref}G_{\rm ref},$ (15)
the absolute Gibbs free energy can be obtained as
$G_{\rm ref}=F^{0}_{\rm ref}+\Delta F^{0}_{\rm ref}+P_{\rm ref}V_{\rm ref}$ (17)
3 Conditional framework
3.1 Conditioning Boltzmann Generators on temperature and pressure
Figure 1: Workflow conditional Boltzmann Generator. The prior equilibrium distribution $q_{{\bf c}_{0}}$ at a reference thermodynamic state ${\bf c}_{0}=(T_{0},P_{0})$ is sampled with MD. A BG conditioned on the thermodynamic variables transforms this distribution to approximate the target equilibrium distributions $q_{{\bf c}_{i}}$ at different thermodynamic states ${\bf c}_{i}$. The BG is trained by minimizing the conditional loss function $\mathcal{L(\theta)}$, given as expectation value over the conditional value ${\bf c}\sim p_{\bf c}$ of the single-point KL divergences (Eq. (19).
The key idea behind our approach is to compute free energy differences over a wide range of thermodynamic states without the need to run MD simulations at each state point. This is achieved by training a BG conditioned on temperature and pressure, which learns the mapping from a prior distribution at a reference thermodynamic state to a whole family of target distributions across thermodynamic space.
Figure 1 illustrates the workflow of our approach. We first sample the prior distribution at the reference state in the $NPT$ ensemble with MD and save samples consisting of the atomic configuration ${\bf x}$ and the box volume $V$. In this work, we consider rectangular boxes and only allow for isotropic scaling, but the proposed approach is completely general and can easily be extended to fully triclinic simulation boxes. Samples $({\bf x},V)$ from the prior distribution $q_{{\bf c}_{0}}$ of the reference state ${\bf c}_{0}=(T_{0},P_{0})$ with temperature $T_{0}$ and pressure $P_{0}$ are transformed by the flow to $({\bf x}^{\prime},V^{\prime})$ in order to approximate equilibrium distributions at multiple different thermodynamic conditions. Both transformations are conditioned on temperature and pressure, which is achieved by feeding the thermodynamic state as an additional input to the model during training (see Sec. 3.2[31].
To optimize the flow parameters, we employ conditional energy-based training (Eq. (13)), avoiding the need to sample the target distributions. Writing the conditional $NPT$ target equilibrium distribution as
$q({\bf x},V|{\bf c})\propto e^{-u({\bf x},V|{\bf c})}$ (18)
with the reduced potential defined in Eq. (1) and adjusting the importance weights in Eq. (8) to the conditional setting, this yields the following loss function:
$\mathcal{L}(\theta)$ $=-\mathbb{E}_{(T,P)\sim p_{\bf c}}\mathbb{E}_{({\bf x},V)\sim q_{{\bf c}_{0}}}\bigl{[}\log w({\bf x},V|{\bf c})\bigr{]}$ (19)
$=\mathbb{E}_{(T,P)\sim p_{\bf c}}\mathbb{E}_{({\bf x},V)\sim q_{{\bf c}_{0}}}\bigl{[}-\beta_{0}(U({\bf x})+P_{0}V)+\beta(U({\bf x}^{\prime})+PV^{\prime})$ (20)
$\hskip 110.96556pt-\log|\det J_{f}({\bf x},V|T,P)|\bigr{]}.$
The conditioning states ${\bf c}\!=\!(T,P)$ are drawn from the distribution $p_{\bf c}$, and $({\bf x}^{\prime},V^{\prime})$ is the sample generated by the conditional flow $f({\bf x},V|T,P)$. After training, equilibrium expectation values of an observable $O({\bf x},V)$ can be computed as
$\mathbb{E}_{({\bf x},V)\sim q_{\bf c}}\,\left[O({\bf x},V)\right]\approx\frac{\mathbb{E}_{({\bf x},V)\sim q_{{\bf c}_{0}}}\left[w({\bf x},V|{\bf c})O({\bf x}^{\prime},V^{\prime})\right]}{\mathbb{E}_{({\bf x},V)\sim q_{{\bf c}_{0}}}\left[w({\bf x},V|{\bf c})\right]}.$ (21)
@@ 3.1 Conditioning Boltzmann Generators on temperature and pressure
The free energy difference between the prior distribution and the target distribution is obtained from TFEP (Eq. (7)) as
@@
Compared to temperature steerable flows [36], the proposed workflow is more general and allows for more flexible models, as no requirements need to be placed on the architecture or the prior.
@@ 3.2 Flow architecture
Figure 2: Flow transformation. a, Schematic of the conditional transformation of one sample $f({\bf x},V)=({\bf x}^{\prime},V^{\prime})$ consisting of an atomic configuration ${\bf x}$ and a volume parameter $V$. The transformation of the box parameter is a simple isotropic scaling operation ($L^{\prime}_{i}/L^{\prime}_{j}\!=\!L_{i}/L_{j}$) learned by an MLP conditioned on $(T,P)$. The configuration ${\bf x}$ is scaled to fractional coordinates ${\bf s}$ and then transformed by a configurational coupling flow $f_{\bf s}$ conditioned on $T$, $P$, and $V^{\prime}$. As a final step, the configuration is scaled back to real coordinates. b, Illustration of the configurational flow transformation, where each flow layer transforms a subset of cartesian coordinates of all particles. The dashed box is a sketch of one $(T,P,V^{\prime})$- conditional coupling layer. The sets of fractional coordinates and spline parameters are denoted as $\{s\}=\{s_{1},\dots,s_{N}\}$ and $\{\psi\}=\{\psi_{1},\dots,\psi_{N}\}$, respectively (see Sec. 3.3 for details). Dashed lines denote splitting / merging operations.
The flow transforms a sample of the prior distribution $({\bf x},V)$ to $({\bf x}^{\prime},V^{\prime})$, where the volume is connected to the box dimensions ${\bf L}$ by $V\!=\!\Pi_{i=1}^{3}L_{i}$. In this work, we focus on the simulation of solids and liquids and therefore apply certain restrictions regarding the shape of the simulation box. We use orthorhombic simulation boxes with ${\bf L}=[L_{1},L_{2},L_{3}]$ and further only allow for an isotropic scaling of the box ($L_{i}/L_{j}\!=\!{\rm const.}$), such that the box is fully characterized by its volume $V\!=\!\Pi_{i=1}^{3}L_{i}$.
Our flow is based on a shape-conditional architecture (see Fig. 2 a[21]. Here, the volume is updated by an affine transformation, while the atomic configurations ${\bf x}\in\mathbb{R}^{3N}$ are first scaled to fractional coordinates ${\bf s}\in[0,1]^{3N}$ and then updated conditioned on the volume and the thermodynamic state. The volume transformation and the transformation of the position of particle $i$ can be written as
@@ 3.3 Conditional coupling layer
As condensed phase systems can exist in the liquid as well as in the solid phase, it is crucial to treat both phases on the same footing. Concretely, $f_{\bf s}$ must respect periodic boundary conditions and invariance under particle permutation. In addition, the center of mass needs to be fixed to prevent uncontrolled errors in the free energy prediction due to rigid translations [20].
To achieve these requirements, we extend the coupling layer originally proposed in [18] and refined in [19] to the case of conditional sampling. Figure 2 b shows our implementation of a $(T,P,V)$ conditional coupling layer. The $N$-particle configuration ${\bf s}$ is split over cartesian coordinates into two subsets $\{s^{\nu}\}\!=\!\{s^{\nu}_{1},...,s^{\nu}_{N}\}$ and $\{s^{\mu}\}\!=\!\{s^{\mu}_{1},...,s^{\mu}_{N}\}$, e.g. $\nu\!=\!\!\{x\}$ and $\mu\!=\!\{y,z\}$. All coordinates indexed by $\nu$ remain unchanged, ${s^{\prime}_{i}}^{\nu}={s_{i}}^{\nu}$ , while the coordinates of atom $i$ carrying index $\mu$ are transformed according to
@@
Here, $G$ is a circular spline [37], which naturally accounts for the periodicity of the problem. The parameters of the spline $\psi_{i}^{\mu}$ are obtained as the output of the network $C$ for the coordinates indexed by $\mu$ of particle $i$, which takes as input circular encodings of the coordinates $\nu$ concatenated with $T$,$P$, and $V^{\prime}$. Following Refs. 18, 19, $C$ is built as a specified number of transformer blocks each consisting of a multi-headed attention [38] layer followed by an MLP. Both layers are implemented as residual updates. The partitioning of $\nu$ and $\mu$ is changed between layers such that each coordinate gets transformed (see [18, 19] for details).
@@
4 Free energy differences and phase diagram
Figure 3: Training results for 180 Lennard-Jones particles in FCC phase. a, KL divergence (Eq. (10)) between generated and target distributions at the evaluation state ${\bf c}_{\rm eval}$ obtained from a flow trained only on the evaluation state (orange) and from a flow trained in a conditional way (blue). Thermodynamic states ($T^{*}_{0},P^{*}_{0}$) and ($T^{*}_{\rm eval},P^{*}_{\rm eval}$) are marked as blue cross and blue star, respectively, in b. Learning rate reductions were applied for the conditional flow after 30 and 45 epochs. The orange dashed lines marks the training time after which the sampling efficiency of the single-point flow was evaluated. b and c, Sampling efficiency (Eq. (9)) for conditional and single-point flow, respectively, in % evaluated over the range of thermodynamic states used for training the conditional flow. d, Deviation in $\Delta G^{*}$ (relative to ($T^{*}_{0},P^{*}_{0}$)) of the conditional flow from MD+MBAR along the path defined by the black arrow in b. The dashed lines indicate one standard deviation evaluated over 10 flow runs, the shaded area correspond to the maximum standard deviation of 10 MD+MBAR runs. e and f, Radial distribution function of configurations from the prior, generated, and reweighted distributions of the conditional flow compared to MD at different thermodynamic states. g, 2D projections of 500 different configurations and mean box dimensions. The left-most panel contains samples from the prior distribution, the remaining panels show configurations generated for increasing temperature and decreasing pressure. For reference, prior configurations and mean box are superimposed to the generated configurations as orange shadows and dashed lines, respectively. In all plots, the fixed particle is located at (0,0).
To measure the performances of the two flow types over training time, we evaluate the KL divergence (Eq. (10), with $\Delta f$ from standard MD+MBAR simulations) between the generated distribution and the target distribution at the evaluation point (Fig. 3 a). The KL divergence of the single-point BG quickly converges within a few epochs, after which an increase in divergence indicates the onset of overfitting. The KL divergence of the conditional BG is much more noisy in the first 20 epochs of training, with a temporary increase before continuing to decrease. The convergence behavior can be explained by the fact that the conditional BG requires more training steps as it learns a whole family of mappings simultaneously. After applying a learning rate reduction, $D_{\text{KL}}$ becomes less noisy and converges to a similar value as the single-point BG. Even though training the conditional BG until convergence requires approximately 10 times more epochs than training the single-point BG, the additional training cost is amortized by the wide range of thermodynamic states covered by the conditional BG.
To assess the accuracy of the flow as free energy estimator, using 2 000 test samples, we compute the Gibbs free energy difference $\Delta G^{*}$ (Eq. (22)) with respect to the prior along the path $P^{*}\in[4,20]$ for a fixed $T^{*}=1.0$ and evaluate the error compared to results obtained with MBAR used across a set of MD simulations (Fig. 3 d). Overall, we observe that both methods are in excellent agreement over the entire range. The accuracy only decreases slightly with increasing distance from the prior and decreasing sampling efficiency. Nevertheless, even for the furthest point, $P^{*}=4$, the mean deviation per particle is only $5\cdot 10^{-3}$, while the absolute value in $\Delta G^{*}$ between $P^{*}=4$ and $P^{*}=20$ is around 14.5 per particle, that is the maximum relative error along the line is less than 0.05%. This demonstrates that the conditional BG can be used for the efficient evaluation of free energy differences also for states far away from the prior.
Furthermore, we illustrate the ability of the model to accurately capture pairwise correlations over varying thermodynamic conditions by computing the radial distribution function (RDF) $g(r)$ [5] for configurations at the start and end point of the path (Fig. 3 e and f) and compare to MD. For both $(T,P)$-conditions, the RDFs resulting from samples of the generated distributions are nearly indistinguishable from the MD results, even before reweighting, while samples from the prior distribution clearly yield different RDFs. Any remaining deviations between the mapped and MD results are resolved by reweighting the generated samples (Eq. (21)). The ability of the conditional BG model to generate samples with varying displacements and volumes is further illustrated in Fig. 3 g, where 2D-projections of 500 generated configurations with increasing temperature and decreasing pressure are shown.
4.2 Mapping out coexistence lines
Figure 4: Coexistence line a, Contourplot of $\Delta G^{*}=G^{*}_{\rm sol}-G^{*}_{\rm liq}$ per particle as obtained from the conditional BGs. Liquid-solid coexistence corresponds to $\Delta G^{*}\!=\!0$. The dashed line denotes the coexistence line obtained from a high-accuracy MD+MBAR run. The locations of solid and liquid priors are marked as blue and red circles, respectively. b, Deviation in the melting temperature of the conditional BG from MD+MBAR. The uncertainties of the BG predictions were obtained as the standard deviation over 10 independent runs. The shaded area indicates the maximum fluctuations over 10 MD+MBAR runs.
Having demonstrated the efficiency of the conditional BG for accurately computing free energy differences over a wide range of thermodynamic states, we now apply this methodology to the exploration of the phase diagram of the LJ system.
Using the LJ system as an example, we demonstrate that our framework based on $(T,P)$-conditioned flows can alleviate these issues. Specifically, we compute the coexistence line between the solid FCC and liquid phase. The only user-defined input to our workflow consists in choosing appropriate locations of the prior states for the two phases of interest. If a point of coexistence is already known, this point is a good choice as the sampling efficiency is highest close to the prior location (see Fig. 3) which allows to compute the coexistence line with only a small number of samples. To emphasize that our approach works without explicit knowledge of the coexistence region, we define a broad range of thermodynamic states with $T^{*}\in[0.6,1.6]$ and $P^{*}\in[1,20]$ in which the coexistence line is expected [41]. Based on physical intuition, we place the prior for the solid at low temperature and high pressure $(T^{*}_{\rm sol}=1,P^{*}_{\rm sol}=14)$, and for the liquid at high temperature and low pressure $(T^{*}_{\rm liq}=1.4,P^{*}_{\rm liq}=6)$. Conditional flows for both phases are trained over the expected region of coexistence (details are given in the SI). After training, we define a grid of points in the $(T^{*},P^{*})$ diagram with a spacing of $(\Delta T^{*},\Delta P^{*})$=(0.05,1.4) and use the BGs to evaluate the relative free energies to the respective prior based on 2000 samples.
In Fig. 4 a, the free energy difference between the solid and liquid phase around the coexistence line is shown as a function of temperature and pressure (see SI for free energy differences over the entire $(T,P)$-range considered during training). The coexistence line predicted by the BG is in excellent agreement with the reference values, obtained from high-accuracy MD+MBAR simulations (dashed line in Fig. 4 a), with only small deviations in the high-temperature region. In Fig. 4 b, the error in the melting temperatures $T^{*}_{\rm melt}$ as a function of pressure are shown as determined by the BG in comparison to MD+MBAR. For the low-pressure regions, the melting temperatures predicted by the BG are well within the error bars of the reference calculations. Only for higher pressures, deviations become slightly larger. Overall, the point of equal free energy, and therefore the melting temperatures, can be determined with a maximum (mean) deviation accuracy of less than $0.01\,T^{*}$.
4.3 Computational efficiency
[…]
5 Discussion
The introduced Boltzmann Generator framework based on conditional normalizing flows provides a novel approach to efficiently calculate differences in free energies over a wide range of thermodynamic states. Contrary to traditional approaches, such as thermodynamic integration or expanded ensembles, it does not require an explicit discretization of the temperature and pressure space. For both solid and liquid phases, flows trained in a conditional fashion achieve large sampling efficiencies even far away from the prior states. As a consequence, only minimal prior knowledge about the system and the melting region is required to map out coexistence lines between two phases with high accuracy. Our workflow, therefore, constitutes a first step towards the efficient and simple calculation of phase diagrams using normalizing flows.